OpenStudy (anonymous):

help? For y = x2 − 4x + 3, Determine if the parabola opens up or down. State if the vertex will be a maximum or minimum. Find the vertex. Find the x-intercepts. Describe the graph of the equation. Show all work and use complete sentences

OpenStudy (anonymous):

the leading coefficient is positive (it is one) so it opens up

OpenStudy (anonymous):

that means the second coordinate of the vertex will be a minimum

OpenStudy (anonymous):

first coordinate of the vertex of \(ax^2+bx+c\) is \(-\frac{b}{2a}\) which in your case is \[-\frac{-4}{2\times 1}=2\] second coordinate is what you get for y when you replace \(x\) by \(2\)

OpenStudy (anonymous):

x intercepts solve \[ x^2 − 4x + 3=0\] which is easy in this case because it factors

OpenStudy (anonymous):

i will leave the complete sentence nonsense up to you i barely speak inglish but here is a nice picture http://www.wolframalpha.com/input/?i=x^2-4x%2B3